# Registration by Regression (RbR): a framework for interpretable and flexible atlas registration

### v1

@@ Abstract
In human neuroimaging studies, atlas registration enables mapping MRI scans to a common coordinate frame, which is necessary to aggregate data from multiple subjects. Machine learning registration methods have achieved excellent speed and accuracy but lack interpretability. More recently, keypoint-based methods have been proposed to tackle this issue, but their accuracy is still subpar, particularly when fitting nonlinear transforms. Here we propose Registration by Regression (RbR), a novel atlas registration framework that is highly robust and flexible, conceptually simple, and can be trained with cheaply obtained data. RbR predicts the $(x,y,z)$ atlas coordinates for every voxel of the input scan (i.e., every voxel is a keypoint), and then uses closed-form expressions to quickly fit transforms using a wide array of possible deformation models, including affine and nonlinear (e.g., Bspline, Demons, invertible diffeomorphic models, etc.). Robustness is provided by the large number of voxels informing the registration and can be further increased by robust estimators like RANSAC. Experiments on independent public datasets show that RbR yields more accurate registration than competing keypoint approaches, while providing full control of the deformation model.
Keywords:
Registration atlas brain ML interpretability
@@ 1 Introduction
Image registration seeks to find a spatial mapping between two images, and is a crucial component of human neuroimaging pipelines. Image registration enables: measuring change between different timepoints of the same subject [17]; building population atlases or subject-specific templates via co-registration of multiple scans [25]; measuring structural differences via tensor-based morphometry [20]; mapping pre-, intra-, and post-operative images [1]; automated segmentation [22]; and many others.
One application of these methods is registering brain MRI scans to digital atlases. Atlases are prototypical brain images that are representative of a population and are often built by co-registering and averaging scans from a cohort of subjects (e.g., the ubiquitous MNI atlas). Atlas registration enables aggregation of spatial information of multiple subjects into a common coordinate frame (CCF). Such CCFs enable studies like tensor- or voxel-based morphometry [5]. Moreover, it also enables the mapping of population-wide information to the native space of a specific subject, e.g., label priors for image segmentation [12]. Atlas registration typically relies on the same methods as pairwise methods.
@@
Most classical registration methods rely on numerical optimization [31] to minimize a cost function, often combining an image similarity term with a regularizer; the former seeks to align the images as close as possible, while the latter prevents excessively convoluted (unrealistic) deformation. Decades of research have provided medical image registration with sophisticated features, e.g., diffeomorphisms [6], symmetry [9], or inter-modality mapping [32].
@@
In 2017, faster registration methods based on deep learning were initially published. The first attempts were supervised and sought to predict ground truth fields obtained, e.g., with accurate classical methods [39]. These algorithms then evolved into unsupervised methods, trained with the same losses used by classical algorithms – possibly augmented with segmentation losses [7, 10]. By now, deep learning registration methods have incorporated many of the features of their classical counterparts (e.g., diffeomorphisms [26], symmetry [21], progressive warping [28], or inter-modality support [14]) and have also started to rely on more modern architectures, e.g., transformers [8] rather than U-nets [7].
@@
Very recently, keypoint methods [37, 15] have been proposed, which predict blob-like feature maps such that their centers of gravity can be used to analytically compute a transform between the two images. This approach addresses two limitations of machine learning registration methods. First, the lack of interpretability associated with the “black box” nature of neural networks, by providing insight into the factors driving the alignment (i.e., the keypoints). And second, the lack of robustness of the affine component. Earlier learning approaches assumed that the inputs were already affinely aligned; some tried to predict the 12 parameter registration matrix [10, 30], but these predictions were generally not robust (e.g., sensitive to initialization) and often failed to generalize well to new datasets. Instead, the affine matrix that best aligns two point clouds can be analytically estimated from keypoints in a robust fashion. Moreover, keypoints can also be used to fit nonlinear transforms, e.g., thin-plate splines (TPS) in KeyMorph [37]).
@@
Here we present Registration by Regression (RbR), a novel framework that enables flexible, robust registration of brain MRI scans to digital atlases. In short, RbR is an evolution of keypoint methods where every voxel is a keypoint: it uses a convolutional neural network (CNN) that predicts, for each voxel of the input scan, its $(x,y,z)$ atlas coordinates. Ground truth can be obtained cheaply via supervision with existing algorithms. This paradigm has a number of advantages: (i) Since the density of keypoints is much higher (over a million vs $\sim$500), nonlinear transforms can be fitted more accurately. (ii) It does not need pretraining to avoid clustering of keypoints in the center of the image. (ii) After a forward pass of the CNN, one can fit with minimal computation (closed-form expressions) a wide array of popular transforms and regularizers, including affine and nonlinear models like Bsplines [34], Demons [35], or the diffeomorphic log-polyaffine framework [2], among others. (iv) Every hyperparameter of the deformation model can be readily modified. This is in contrast with hypernetwork approaches [18], where only specific parameters can be tuned as the space of deformations is constant (e.g., control point spacing cannot be specified). (v) It can achieve enhanced robustness using robust estimators like RANSAC [13]. (vi) It is easy to implement, and ground truth for a new atlas or dataset can be easily obtained with classical registration methods.
@@ 2 Methods
Our proposed method is summarized in Figure 1. It comprises three different stages: data preparation, training, and test-time fitting.
@@ 2.1 Data preparation
The crux of RbR is a regression CNN that estimates, for every voxel of the input scan (with discrete coordinates $x,y,z$), its corresponding atlas coordinates $x^{\prime},y^{\prime},z^{\prime}$. This CNN is trained in a supervised fashion using accurate nonlinear deformations obtained with a classical registration approach $\mathcal{R}$. If $I(x,y,z)$ is the input image and $A$ is a reference atlas, the ground truth registration is:
@@
Therefore, a large labeled dataset can be obtained “for free” by simply registering a set of unlabeled images to the atlas at hand using a well-established registration algorithm. While there are no formal requisites on the deformation model of the registration method, diffeomorphic models that yield smooth and invertible fields are preferred [6]. Unlabeled images exist in abundance in public datasets – particularly 1mm isotropic MPRAGEs typically used in neuroimaging research.
Figure 1: (a) Training data are prepared by nonlinearly registering T1w scans from HCP and ADNI to the MNI template using NiftyReg. (b) A U-Net is trained to predict the MNI coordinates for every voxel of an input scan. (c) At test time, almost any transformation can be fitted between the input scan and predicted MNI coordinates.
@@ 2.2 Training
Training RbR is simple. Given a 3D voxel-wise regression CNN $\bm{N}(\cdot)$ with parameters $\bm{\theta}$, an atlas $A$, and a set of $J$ images $\{I_{j}\}_{j=1,\ldots,J}$, the goal is to minimize the following loss function with stochastic gradient descent:
@@
where $\Omega_{j}$ is the image domain of image $I_{j}$ and is typically constrained to brain regions (i.e., excludes extracerebral tissue). Equation 1 is just the $\ell_{1}$ norm between the predicted and ground truth atlas coordinates; we use $\ell_{1}$ rather than $\ell_{2}$ due to its higher robustness. We note that no regularization is needed: the CNN just does its best in a voxel-by-voxel basis, smoothness is provided by the deformation model and its parameters at test time. However, as with any other CNN, augmentation is crucial to endow the trained models with the best possible generalization ability. We use aggressive augmentation during training, including random blurring, bias field, noise, and deformation (affine and elastic).
@@ 2.3 Test-time fitting
The flexibility and robustness of RbR stem from its test-time fitting strategy. Given an input image $I$, one first processes it with the trained neural network to obtain $[x^{\prime},y^{\prime},z^{\prime}]=\bm{N}(x,y,z;I,\hat{\bm{\theta}})$; we note that the atlas is not required at this point. Then, one can fit a deformation model to this predicted field to obtain the final atlas registration. We consider three different types of models in this article: a family of models based on basis functions, a model based on the Demons algorithm [35], and a diffeomorphic model based on the log-polyaffine framework [2]; we note that other models are possible.
Basis functions: a large class of transforms that can be written as a sum of spatial basis functions can be fitted to $[x^{\prime},y^{\prime},z^{\prime}]$ using regularized least squares, one spatial coordinate at the time. Let $\bm{d}$ be the flattened column vector of $x^{\prime}$, $y^{\prime}$, or $z^{\prime}$ coordinates. Let $\bm{\phi}$ be a set of $B$ spatial functions evaluated on the discrete image domain $\Omega$, such that each row corresponds to a spatial location (flattened), and each column corresponds to a basis function. Finally, let $\bm{c}$ be a column vector with $B$ elements, corresponding to the coefficients of the basis functions. Then, the fitted transform is $\bm{\phi}\bm{c}$ and its squared coordinate error is:
@@
Other basis previously used in registration that are straightforward to implement include lower-order polynomials [38] and the discrete cosine transform [4].
@@
Log-polyaffine model: Diffeomorphic models that can be analytically inverted are desirable in registration. Here we explore a log-polyaffine model [2] that is fitted to the output of the CNN as follows. First, we subdivide the image domain into cubic supervoxels of fixed, user-defined width $W$; this parameter controls the flexibility of the transform. Next, we compute an affine transform for every supervoxel $s$, assuming that it contains a minimum number of voxels. If we denote this affine transform by $\bm{T}_{s}$ in homogeneous coordinates, then [2]:
@@ 2.4 Implementation details
The CNN is a standard U-net [33], with a design inspired by nnU-net [23]. It has four resolution levels with two convolutional layers (comprising $3\times 3\times 3$ convolutions and a ReLu) followed by $2\times 2\times 2$ max pooling (in the encoder) or upconvolution (decoder). The final activation layer is linear, to regress the atlas coordinates in decimeters (which roughly normalizes them from -1 to 1). In addition to predicting atlas coordinates, the CNN also learns to predict a brain mask, which is used to define the domain $\Omega$ at test time. For this, we add a segmentation loss to Equation 1 with relative weight 0.5; the segmentation loss itself combines a Dice loss (weight 0.75) and cross-entropy loss (weight 0.25). The model is trained with stochastic gradient descent with learning rate 0.01, weight decay $3\mathrm{e}{-5}$, and momentum 0.99 for 100 epochs, setting aside 20% of the data for validation (selecting the best model). Input images are normalized with their median value inside the brain mask. At test time, nonlinear models are fitted on the residual of the Affine-RANSAC fit, clipped at 10mm to minimize the impact of the worst outliers. The RANSAC algorithm uses 50,000 randomly selected voxels, a maximum of 100 iterations, tests 500 voxels in every attempt, and requires 20% of voxels to be inliers to consider any given solution. The minimum number of voxels per supervoxel in the log-polyaffine model is 100.
@@ 3.2 Experimental setup
We quantify registration accuracy with average Dice scores between the segmentations of the scans and the segmentation of the atlas, deformed with the estimated registrations. We further evaluate the regularity of the deformations using the membrane energy. Using these metrics, we compared RbR with:
NiftyReg: We use the same parameters as for the training data, Since we used it as ground truth in training, NiftyReg provides a ceiling for the performance that RbR can achieve.
@@
Figure 3: Left: Average Dice between segmentations of scans and registered MNI. For RbR, the numbers indicate control point spacing (Bsplines), Gaussian $\sigma$ (demons), or supervoxel width (log-polyaffine). Right: membrane energy for nonlinear models.
@@ 3.3 Results
Figure 2 shows qualitative results for the different methods, while Figure 3 shows box plots for the average Dice and membrane energy. KeyMorph with TPS yields an average Dice of 0.60, which is 7 points lower than the value reported in the original publication. This could be a domain gap issue – including both the test datasets and the MNI atlas, which is blurrier than a regular scan. Errors are apparent in Figure 2, where KeyMorph fails to correct the rotation about the A-P axis in linear mode and introduces wrong deformations in nonlinear mode.
@@ 3.4 Discussion and conclusion
RbR offers a new perspective on interpretable keypoint-based registration, posing it as a coordinate regression problem. Allowing every voxel to inform the model fit enables RbR to outperform standard keypoint approaches and fit less parsimonious nonlinear models. RbR also has disadvantages, particularly the need to retrain the model for each new atlas. Future work will include fitting other models, adding topological losses in training (e.g., penalizing negative Jacobians), investigating RbR’s value as a feature extractor (i.e., as pretraining for other tasks), and exploring other improvements that help close the gap with non-interpretable approaches. We believe that RbR will be particularly valuable in scenarios requiring robust and interpretable registration, such as fetal imaging.
Acknowledgments
Work primarily supported by NIH grants 1R01AG070988 and 1UM1MH130981. Support also provided by NIH 1RF1MH123195, 1RF1AG080371, and R00-HD101553. Oula Puonti was supported by a grant from Lundbeckfonden (grant number R360–2021–395).

### v2

@@ Abstract
In human neuroimaging studies, atlas registration enables mapping MRI scans to a common coordinate frame, which is necessary to aggregate data from multiple subjects. Machine learning registration methods have achieved excellent speed and accuracy but lack interpretability and flexibility at test time (since their deformation model is fixed). More recently, keypoint-based methods have been proposed to tackle these issues, but their accuracy is still subpar, particularly when fitting nonlinear transforms. Here we propose Registration by Regression (RbR), a novel atlas registration framework that: is highly robust and flexible; can be trained with cheaply obtained data; and operates on a single channel, such that it can also be used as pretraining for other tasks. RbR predicts the $(x,y,z)$ atlas coordinates for every voxel of the input scan (i.e., every voxel is a keypoint), and then uses closed-form expressions to quickly fit transforms using a wide array of possible deformation models, including affine and nonlinear (e.g., Bspline, Demons, invertible diffeomorphic models, etc.). Robustness is provided by the large number of voxels informing the registration and can be further increased by robust estimators like RANSAC. Experiments on independent public datasets show that RbR yields more accurate registration than competing keypoint approaches, over a wide range of deformation models.
@@ 1 Introduction
Image registration seeks to find a spatial mapping between two images, and is a crucial component of human neuroimaging pipelines, e.g., for measuring change between different timepoints of the same subject [17]; building population atlases or subject-specific templates via co-registration of multiple scans [25]; measuring structural differences via tensor-based morphometry [20]; mapping pre-, intra-, and post-operative images [1]; or automated segmentation [22]. One important application of these methods is registering brain MRI scans to digital atlases, such as MNI. Atlas registration enables aggregation of spatial information of multiple subjects into a common coordinate frame (e.g., for voxel-based morphometry [5]), as well as mapping of population-wide information to the native space of a specific subject, e.g., label priors for image segmentation [12]. Atlas registration typically relies on the same methods as pairwise registration.
@@
Most classical pairwise registration methods use numerical optimization [31] to minimize a cost function, often combining an image similarity term with a regularizer; the former seeks to align the images to each other, while the latter prevents excessively convoluted (unrealistic) deformation. Decades of research have provided medical image registration with features such as diffeomorphisms [6], symmetry [9], or inter-modality mapping [32].
@@
In 2017, faster registration methods based on deep learning were initially published. The first attempts were supervised and sought to predict ground truth fields obtained, e.g., with accurate classical methods [39]. These algorithms then evolved into unsupervised methods, trained with the same losses used by classical algorithms – possibly augmented with segmentation losses [7, 10]. By now, deep learning registration methods have incorporated many of the features of their classical counterparts, e.g., diffeomorphisms [26], symmetry [21], progressive warping [28], or inter-modality support [14]. However, deep learning registration also has drawbacks, including: lack of interpretability (as many other deep learning domains); lack of flexibility of the deformation model, which is fixed and cannot be modified at test time (beyond tweaking the regularizer weight with hypernetworks [18]); and lack of robustness of the affine component.
@@
Very recently, keypoint methods [37, 15] have been proposed, which predict blob-like feature maps such that their centers of gravity can be used to analytically compute a transform between the two images. While the feature prediction remains opaque, this approach partly address the lack of interpretability by providing insight into the factors driving the alignment, i.e., the keypoints. Keypoints also have potential in dynamic selection of deformation model at test time to warp one point cloud to the other, e.g., affine or thin-plate spline (TPS) transforms in KeyMorph [37]. Finally, keypoints also effectively tackle the lack of robustness of the affine component. Earlier learning approaches for nonlinear registration assumed that the inputs were already affinely aligned; some tried to predict the 12-parameter affine matrix [10, 30], but these predictions were generally sensitive to initialization and often failed to generalize well to new datasets.
@@
Here we present Registration by Regression (RbR), a novel framework for flexible, robust registration of brain MRI scans to atlases. RbR is an evolution of keypoint methods where every voxel is a keypoint: it is a supervised convolutional neural network (CNN) that predicts $(x,y,z)$ atlas coordinates for every input voxel. RbR has a number of advantages: (i) Since the number of keypoints is very high ($\sim 10^{6}$ vs $\sim 10^{2}$ in KeyMorph), nonlinear transforms can be fitted more accurately. (ii) After a forward pass of the CNN, a wide array of popular transforms and regularizers can be fit in closed-form, including affine and nonlinear models like Bsplines [34], Demons [35], or the diffeomorphic log-polyaffine model [2], among others. (iii) It supports robust estimators like RANSAC [13]. (iv) Every hyperparameter of the deformation model can be readily modified; this is in contrast with hypernetwork approaches [18], where only specific parameters can be tuned as the space of deformations is constant (e.g., control point spacing cannot be specified). (v) Being a prediction on a single image (channel), it can be directly and easily used as pretraining (“feature extractor”) for other tasks. (vi) Compared with KeyMorph, it does not need pretraining to avoid clustering of keypoints in the center of the image. (vii) Compared with classical registration, it is much faster (particularly when considering that one fit different transforms “for free” after the forward pass) and can be used as building block in deep learning architectures (as a frozen differentiable block to align images, or as a feature extractor as mentioned above).
@@ 2.1 Data preparation
RbR (Figure 1) is a regression CNN that estimates, for every voxel of the input scan (with discrete coordinates $x,y,z$), its corresponding atlas coordinates $x^{\prime},y^{\prime},z^{\prime}$. RbR uses supervised training with accurate nonlinear deformations obtained with a classical registration approach $\mathcal{R}$. If $I(x,y,z)$ is the input image and $A$ is an atlas, the ground truth registration is:
@@
Therefore, a large labeled dataset can be obtained “for free” by simply registering a set of unlabeled images to the atlas at hand using a classical registration algorithm – ideally a diffeomorphic methods that yields smooth and invertible fields  [6]. We note that unlabeled images exist in abundance in public datasets – particularly 1mm isotropic MPRAGEs typically used in neuroimaging research.
We also note that, compared with using synthetic fields, the performance of our strategy is bounded by that of the classical algorithm $\mathcal{R}$. However, synthesis of realistic (atlas, image, deformation) triplets is not trivial, since atlases are much smoother than images. Moreover, classical algorithms still provide state-of-the-art accuracy in atlas registration – speed limitations aside.
@@ 2.2 Training
Given a 3D voxel-wise regression CNN $\bm{N}(\cdot)$ with parameters $\bm{\theta}$, an atlas $A$, and a set of $J$ images $\{I_{j}\}_{j=1,\ldots,J}$, the goal is to minimize the loss:
@@
where $\Omega_{j}$ is the foreground of $I_{j}$ (i.e., voxels inside the brain), and we use $\ell_{1}$ rather than $\ell_{2}$ due to its higher robustness. We note that no spatial regularization is needed: the CNN just does its best in a voxel-by-voxel basis, and smoothness is provided by the deformation model and its parameters at test time. Nevertheless, augmentation is crucial to endow the trained models with the best possible generalization ability. We use aggressive augmentation during training, including random blurring, bias field, noise, and deformation (affine and elastic).
Figure 1: (a) Training data are prepared by nonlinearly registering T1w scans from HCP and ADNI to the MNI template using NiftyReg. (b) A U-Net is trained to predict the MNI coordinates for every voxel of an input scan. (c) At test time, a wide array of transformations can be fitted between the input scan and predicted MNI coordinates.
@@ 2.3 Test-time fitting
Given a CNN prediction $[x^{\prime},y^{\prime},z^{\prime}]=\bm{N}(x,y,z;I,\hat{\bm{\theta}})$ (which does not require the atlas), one can fit a deformation model to $[x^{\prime},y^{\prime},z^{\prime}]$ to obtain the final atlas registration. We consider three different types of models in this article: a family of models based on basis functions, a model based on the Demons algorithm [35], and a diffeomorphic model based on the log-polyaffine framework [2].
Basis functions: a large class of transforms can be written as a sum of spatial basis functions and can be fitted to $[x^{\prime},y^{\prime},z^{\prime}]$ using regularized least squares, one spatial coordinate at the time. Let $\bm{d}$ be the flattened column vector of $x^{\prime}$, $y^{\prime}$, or $z^{\prime}$ coordinates. Let $\bm{\phi}$ be a set of $B$ spatial functions evaluated on the discrete image domain $\Omega$, such that each row corresponds to a spatial location (flattened), and each column corresponds to a basis function. Finally, let $\bm{c}$ be a column vector with $B$ elements, corresponding to the coefficients of the basis functions. Then, the fitted transform is $\bm{\phi}\bm{c}$ and its squared coordinate error is:
@@
Other basis previously used in registration that would be straightforward to implement include lower-order polynomials [38] or the discrete cosine transform [4].
@@
Log-polyaffine model: Diffeomorphic models that can be analytically inverted are desirable in registration. Here we explore a log-polyaffine model [2] that is fitted to the output of the CNN as follows. First, we subdivide the image domain into cubic supervoxels of fixed, user-defined width $W$; this parameter controls the flexibility of the transform. Next, we compute an affine transform for every supervoxel $s$, assuming that it contains a minimum number of voxels in the brain mask. If we denote this affine transform by $\bm{T}_{s}$ in homogeneous coordinates, if can be shown that [2]:
@@ 2.4 Implementation details
The CNN is a standard U-net [33], with a design inspired by nnU-net [23]. It has four resolution levels with two convolutional layers (comprising $3\times 3\times 3$ convolutions and a ReLu) followed by $2\times 2\times 2$ max pooling (in the encoder) or upconvolution (decoder). The final activation layer is linear, to regress the atlas coordinates in decimeters (which roughly normalizes them from -1 to 1). In addition to predicting atlas coordinates, the CNN also learns to predict a brain mask, which is used to define the domain $\Omega$ (i.e., to mask the loss during training, and to select the voxels that are used to fit the transforms at inference). For this, we add a segmentation loss to Equation 1 with relative weight 0.5; the segmentation loss itself combines a Dice loss (weight 0.75) and cross-entropy loss (weight 0.25). The model is trained with stochastic gradient descent with learning rate 0.01, weight decay $3\mathrm{e}{-5}$, and momentum 0.99 for 100 epochs, setting aside 20% of the data for validation (selecting the best model). Input images are normalized with their median value inside the brain mask. At test time, nonlinear models are fitted on the residual of the Affine-RANSAC fit (clipped at 10mm for outlier rejection). The RANSAC algorithm uses 50,000 randomly selected voxels, a maximum of 100 iterations, tests 500 voxels in every attempt, and requires 20% of voxels to be inliers to consider any given solution. The minimum number of voxels per supervoxel in the log-polyaffine model is 100.
@@ 3.2 Experimental setup
We quantify registration accuracy with average Dice scores between the segmentations of the scans and the segmentation of the atlas, deformed with the estimated registrations. We further evaluate the regularity of the deformations using the membrane energy, which is one of the smoothness metrics used in the registration literature. Using these metrics, we compared RbR with:
NiftyReg: We use the same parameters as for the training data. Since we used it as ground truth in training, NiftyReg provides a ceiling for the performance that RbR can achieve.
@@
Figure 3: (a) Box plot of average Dice between segmentations of scans and registered MNI. For RbR, the numbers indicate control point spacing (Bsplines), Gaussian $\sigma$ (demons), or supervoxel width (log-polyaffine). (b) Significance matrix for two-tailed t-tests comparing the Dice scores; pink represents p$<$0.05. (c) Membrane energy for nonlinear models. In the box plots, the center line is the median, the ends of the box are the 25th and 75th percentiles, and the whiskers extend to the furthest observations not considered outliers – which are marked with circles. Outliers are values more than 1.5 times the interquartile range away from the ends of the box.
@@ 3.3 Results
Figure 2 shows qualitative registration results, while Figure 3 shows box and significance plots for the average Dice and membrane energy. KeyMorph with TPS yields an average Dice of 0.60, which is 7 points lower than the value reported in the original publication. This could be a domain gap issue – including both the test datasets and the MNI atlas, which is blurrier than a regular scan. Errors are apparent in Figure 2, where KeyMorph fails to correct the rotation about the A-P axis in linear mode and introduces wrong deformations in nonlinear mode.
@@ 3.4 Discussion and conclusion
RbR offers a new perspective on interpretable keypoint-based registration, posing it as a coordinate regression problem. Allowing every voxel to inform the model fit enables RbR to outperform standard keypoint approaches and fit less parsimonious nonlinear models. RbR also has disadvantages, particularly the need to retrain the model for each new atlas. Future work will include fitting other models, adding topological losses in training (e.g., penalizing negative Jacobians), investigating RbR’s value as a feature extractor (i.e., as pretraining for other tasks), exploring other improvements that help close the gap with non-interpretable approaches, and integrating with domain randomization approaches for contrast- and resolution-agnostic registration (as in [15]). We believe that RbR will be particularly valuable in scenarios requiring robust and interpretable registration in presence of extreme rotations, such as fetal imaging.